OpenStudy (anonymous):

Given equation A as 2x+1/4y=3 and equation B as 2/3x-y=6, which expression will give the value of y?

OpenStudy (anonymous):

A. a-3b B. a-3/2b C. 2a-3b D. 1/3a+b

OpenStudy (phi):

it looks like they are asking how to combine equation A and equation B to eliminate the x variable eq A is 2x+1/4y=3 eq B is 2/3x-y=6 if you multiply eq B by -3 (that means both sides, and all terms) you get -3*⅔ x + -3*-y = -3*6 that simplifies to -2x + 3y = -18 what do you get if you add that to eq A ?

OpenStudy (anonymous):

3 1/4y=-15

OpenStudy (phi):

and the point is we can solve for y ... write 3 ¼ as 13/4 to get 13/4 y = -15 then multiply both sides by 4/13 to solve for y y = -15*4/13 but we don't need to find y, we just need to answer the question equation A - 3 equation B allows us to solve for y

OpenStudy (anonymous):

thank you

OpenStudy (phi):

yw